OpenStudy (anonymous):

At the store deli, you spent $13.45 on 5 pounds of cheese. How much would it cost to buy 2 pounds of cheese?

OpenStudy (anonymous):

$5.38 ?

OpenStudy (kirbykirby):

\[ \begin{align} 13.45 &\rightarrow 5 \text{ lbs}\\ x &\rightarrow 2 \text{ lbs}\end{align}\] Think of this as a proportion, and then cross-multiply to find x: \[ \frac{13.45}{x}=\frac{5}{2}\] Or another way to think about it is to find the cost per 1 pound, so divide 13.45/5. Then to find the cost for 2 lbs, just multiply by 2